# Applications to Monolayers, Nanotubes and Nanotori

### v1

@@ Abstract
In this work, we present an empirical force field for hexagonal boron nitride (hBN) monolayers, derived via a bottom-up strategy from first principles calculations. We aim to deliver a simple analytical force field for boron nitride which is efficient and applicable to large-scale simulations, without compromising its accuracy. The force field is developed with the goal to analytically reproducing the potential energy contributions arising from planar bond-stretching and bond-angle-bending deformations, as well as from out-of-plane torsional deformations, as obtained from periodic density functional theory calculations. Analytical anharmonic potential energy functions were employed to describe and parameterize the force field through a fitting process. The potential is applied for estimating in-plane (stiffness matrix and elastic constants) and out-of-plane (bending and Gaussian rigidity) mechanical properties of monolayer hBNs using both molecular mechanics calculations and analytical formulas. Additionally, we determined the elastic properties of more complex nanostructures, including hBN nanotubes and nanotori. An analytical expression has been provided for the dependence of the bending rigidity of relatively large diameter hBN nanotubes on their radius.
@@ I Introduction
Hexagonal boron nitride (hBN) has sparked significant interest within the scientific and industrial communities due to its unique properties. Being a two-dimensional material with a wide indirect bandgap (5.955 eV)[1], excellent chemical and thermal stability,[2] exceptional thermal conductivity[3] and good dielectric and optical properties,[4, 5] makes it highly versatile for numerous applications.[6, 2, 5, 7] In the field of electronics, hBN has been successfully employed in transistors, integrated circuits, and capacitors, leading to enhanced performance and improved energy conversion efficiency.[8, 9, 10] Moreover, its remarkable thermal management capabilities have been utilized in electronic devices and green LEDs, effectively dissipating heat.[11] The optical properties of hBN have also proven valuable in optoelectronics and photonics,[12, 13] where it can serve as a transparent electrode and UV light emitter,[14] and as UV-detector.[5] In addition, hBN has been considered for applications in photocatalytic reductions for reducing carbon dioxide.[15] Furthermore, hBN’s chemical stability and lubricating properties have contributed to its use in protective coatings and high-temperature lubrication.[16, 17, 18]
@@
The mechanical properties of hBN have been investigated in the literature via experiments,[23, 24] ab-initio calculations,[25, 26, 27, 28, 29] and atomistic molecular dynamics (MD) simulations using Tersoff-based force fields.[30, 31, 32, 33] In this work, we develop a semi-empirical force field for hBN monolayers encompassing both in-plane (bond-stretching and bond-angle-bending) and out-of-plane (torsional) interactions. The force-field parameterization results from a bottom-up fitting procedure in terms of matching the potential energy of properly deformed hBN geometrical structures as derived from first-principles density functional theory (DFT) calculations.
The functional expressions of the proposed potential are simple yet effective, capable of efficiently addressing very large hBN sheets with large-scale molecular dynamics simulations. The resulting force field is significantly more efficient than previous Tersoff-based implementations[30, 31, 32, 33]. It has been employed to investigate the structural relaxation and elastic properties of various hBN-based nanostructures, including hBN monolayers, nanotubes (NT), and hBN nanotori.
In the following, we provide a detailed description of the parametrization of the in-plane (Section II.1) and out-of-plane (Section II.2) terms of the force field through the fitting of the potential energy of deformed hBN monolayers obtained from ab-initio calculations. Subsequently, in Section III.1 we determine the elastic constants of single-layer (SL) hBNs at T = 0 K, such as the components of stiffness matrix and the corresponding Lamé parameters, Young modulus, Poisson ratio, and bulk modulus. The aforementioned parameters are estimated via molecular mechanics calculations and with analytical closed-form formulas.[34] In Section III.2, we illustrate the variation of the potential energy of hBN nanotubes as a function of their radius and chirality. In the limit of infinite radius, we extrapolate the bending rigidity of the corresponding hBN monolayers to provide comparisons against analytical expressions. Finally, in Section III.3 following a similar procedure for the energy of folded hBN NT in order to form closed nanotori, the bending rigidity of hBN nanotubes is evaluated.
@@ II Force Field Derivation
To simulate the in-plane deformations we considered the minimal 2-atom unit cell and an $\times 8$ Monkhorst pack sampling of the reciprocal space. To determine the bond-stretching potential term, the unit cell lattice vectors were scaled uniformly (Fig. 1a). Regarding the bond-angle-bending term, the unit cell was deformed in such a way (Fig. 1b) that all bond lengths assumed rigid and only the angles were varied in a symmetric fashion (see Fig. 3 below).
For the out-of-plane deformations, we considered deformed BN nanoribbons that were bent around a particular middle axis. Three different cases were considered: (i) bending an 18 atom unit cell around a zig-zag (ZZ) middle axis with only Boron atoms lying on that axis (Fig. 1c top, where the rotational zig-zag axis is shown by the red dotted line denoted as ZZ), (ii) similarly bending a nanoribbon, with a 18 atom unit cell as well, around a zig-zag middle axis passing strictly through Nitrogen atoms, and (iii) bending a nanoribbon, with a 22 atom unit cell, around an armchair (AC) middle axis (Fig. 1c bottom, where the rotational armchair axis is shown by the red dotted line denoted as AC). For the out-of-plane deformations, at least an 8×1 Monkhorst pack sampling of the reciprocal space was used.
@@ II.1 In-plane terms of the force field
Figure 1: Schematic illustration of the deformed hBN structures considered in the DFT calculations. In-plane deformations: (a) affine deformation with a uniform scaling along lattice vectors (Eq. (1)) and (b) deformation altering in-plane angles and maintaining bond lengths (Eq. (4)). Red arrows show the corresponding displacement in each case. Out-of-plane deformations: (c) folding an hBN ribbon by an angle $\varphi$ around a rotation axis (red dotted line) oriented along the ZZ (top) or AC (bottom) direction passing through its plane. In all cases, the dashed lines depict the edges of the deformed unit cell used in the calculation. For clarity, supercells with three unit cell replications along the periodic directions are drawn. The blue transparent (opaque) lines illustrate the periodic boundaries of the supercells before (after) the deformation.
The in-plane potentials of hBN were determined by deforming a periodic primitive cell with the following (unperturbed) lattice vectors: $\mathbf{a}_{1}=(a,0)$ and $\mathbf{a}_{2}=(\frac{a}{2},\frac{a\sqrt{3}}{2})$, where $a=\|\mathbf{a}_{1}\|=\|\mathbf{a}_{2}\|=\sqrt{3}\;l_{\textrm{BN}}$ is the lattice constant and $l_{\textrm{BN}}$ the B-N bond length. In the reference configuration ($\mathbf{R}^{\textrm{ref}}$), where there is no residual stress, the bond length equals $l_{\textrm{BN}}=1.4518$  and the lattice constant is $a=2.5146$.
@@
with $\epsilon$ being the engineering strain and $I_{2}$ the identity matrix with size 2; see illustration in Fig. 1a. This operation scales the bond lengths uniformly and does not modify the in-plane angles. The same method has been applied in Ref. 40 for the case of graphene monolayers.
The bonded interactions of the force field were then parameterized by fitting the DFT energy per bond results obtained from the uniform scaling deformations, shown in Fig. 2 by circles, with the Morse potential
@@
where $D_{\textrm{b}}$ is the potential well depth, $r_{\textrm{b}}$ the equilibrium bond distance, and $a_{\textrm{b}}$ the width parameter of the potential. The fitting of the DFT data with Eq. (3) is shown by the continuous line in Fig. 2, yielding the following parameter values: $D_{\textrm{b}}=5.35\pm 0.01$ eV, $\alpha_{\textrm{b}}=1.897\pm 0.004$-1, and $r_{\textrm{b}}=1.4518$. The calculated equilibrium B-N distance leads to a lattice constant $a=2.5146$  in excellent agreement with literature values [23, 41, 32, 26, 27, 28].
@@
The bond-angle-bending term of the force field is obtained through alteration of the in-plane angles of a pristine hBN, by $-x$ or $x/2$ while maintaining the equilibrium value of $1.4518$ in all B-N bonds, as shown in Fig. 3. [40] In this case, four angles within each hBN ring increase by $x/2$, and two angles decrease by $-x$. This operation deforms the reference configuration nonaffinely. First, we set the origin of the first (1) and second (2) sublattice (e.g., the pink and blue atoms in Fig. 1b, respectively) to
@@ II.2 Out-of-plane terms of the force field
Note that hBN exhibits mirror symmetry with respect to normal planes passing through neighboring B-N atoms; hence, a single deformation suffices for determining folding around the AC direction. The torsional energy of the AC-folded hBNs, $U\mathrm{{}_{t}^{a}}$ is shown in Fig. 6 by black circles. On the contrary, hBN is asymmetric with respect to normal planes passing through second neighboring identical (B or N) atoms. To capture this effect in the case of folding around the ZZ direction, the torsional energy was averaged from two separate folding deformations, where the rotational axis passes through second neighboring either B or N atoms (see Fig. 5 left). The corresponding torsional energies for the two separate ZZ folding deformations are shown by red open symbols in Fig.6, while their average is represented by filled red circles.
Note that the total deformation energy obtained by DFT in these cases includes contributions not only from torsional terms but also from bond-angle-bending terms. Before fitting the parameters of the out-of-plane potentials one must remove the latter contributions, in order to isolate the dependence of the pure torsional terms on the folding angle $\mathit{\phi}$. Details of this fitting procedure can be found in Ref. 43.
Figure 6: Torsional energy of hBN nanoribbons folded along AC (filled black circles) or ZZ (red) rotation axes. In the latter case, the results are shown by empty diamonds when the ZZ axis passes through N atoms and by empty circles when the axis passes through B atoms, as it is illustrated in Fig. 5 (left). The average of these two situations (filled red circles) is considered in the fitting procedure regarding the rotation around ZZ folding axis. Solid lines indicate the model’s fittings using Eqs. (10) and (11) for nanoribbons folded along the AC and ZZ axes.
As it is customary, we define the torsional angle, $\omega$, for a quadruple of atoms $(ijkl)$ as the dihedral angle of the planes defined by $(ijk)$ and $(jkl)$ atom triplets. One can distinguish two kinds of such angles, termed cis and trans. As in ref. 43, we used the following analytic formulas for the out-of-plane distortions as functions of $\mathit{\omega}$:
@@ III.1 In-plane Elastic Constants of hBN Monolayers
where now pair indices are denoted by single numbers, $xx=1$, $yy=2$, and $xy=6$. In doing so, the strain tensor is described by the vector
@@
To determine the elements of the stiffness matrix via MM computations we utilize three fundamental deformation modes (elementary deformations), each characterized by a single non-zero strain component in Eq. 15; hence, the coefficients of the stiffness matrix are derived from the slope of the corresponding stress-strain plots in the linear regime for small strains.[67]
@@
respectively; note that $\epsilon_{1}=\epsilon_{2}=\xi$. The third deformation mode $\bm{\epsilon}=(0,0,\epsilon_{6})$ entails varying the angle $\hat{\gamma}$ between the lattice vectors of an initially orthorhombic box (shearing) upon applying the linear transformer
@@
where $\epsilon_{6}=\tan(\hat{\gamma}\xi)$. The slope of the corresponding stress-strain plots was evaluated with a central-difference scheme with perturbation $\xi=\pm 10^{-7}$. For additional details regarding the determination of the stiffness matrix with explicit deformation, the interested reader is referred to the article by Clavier et al;[67] see Section 2.1 therein. Details regarding the calculation of the stress tensor can be found in Ref. 68.
@@
with $l_{0}$ being the characteristic thickness of the monolayer, $\mathcal{X}$ a quantity with 2D stress units (Pa m) and $\mathcal{X}^{\mathrm{3d}}$ a quantity with 3D stress units (Pa). Care should be exercised when converting between 3D stress and 2D stress values, since various conventions have been invoked across the literature concerning the SL thickness.[32, 31, 24, 29] The results of our calculations in units of 2D stress are illustrated in Table 2.
Table 2:  The lattice parameter, the in-plane and the out-of-plane elastic constants of hBN monolayers, calculated through MM simulations using the proposed force field, along with comparisons of corresponding experimental and theoretical (atomistic or ab initio methods) values reported in the literature1.
@@
$a$ (Å) 2.5146 2.504 [23], 2.505 [41] 2.505 [32] 2.512 [26], 2.51 [27], 2.504 [28]
$C_{11},C_{22}$ (Pa m) 279.309 270 [24] 274.98 [32] 293.2 [26], 289.8 [27]
$C_{12}$ (Pa m) 76.870 56.3 [24] 81.91 [32] 66.1 [26], 63.7 [27]
$C_{66}$ (Pa m) 101.220 - 96.68 [32] 107.92 [29], 113.1 [27]
$\lambda$ (Pa m) 76.870 - - -
$\mu=G$ (Pa m) 101.220 - - -
$\nu$ 0.275 - 0.297 [32] 0.216 [29], 0.2176 [26]
$Y$ (Pa m) 258.153 - 250.5 [32], 293.4 [31], 274.7 [30] 262.3 [29], 278.3 [26], 271[25]
275.8 [27], 256.245 [28]
$B$ (Pa m) 178.089 - - 177 [27]
$D_{\textrm{sheet}}$ (eV) 1.563 - 0.56 [32], 0.86 [69], 0.89 [70] 1.29 [71], 0.95 [28]
$D_{\textrm{G,sheet}}$ (eV) $-3.233$
1 The characteristic thickness ($l_{0}$) for converting 3D stress to 2D units, Eq. (16), has been considered 3.34 Å  in Ref. [32], 3.33 Å  in Refs. [31] and [24], and $r_{\mathrm{N}}+r_{\mathrm{B}}=1.55+1.92=3.47$  in Ref. [29] ($r_{\mathrm{N}}$ and $r_{\mathrm{B}}$ are the van der Waals radii of N and B atoms, respectively).
The coefficients of the stiffness matrix are in good agreement with the values reported in the literature from MD [32] and ab initio [25, 26, 29] simulations. In addition, the lattice constant conforms with the reported experimental [23, 41] and theoretical [32, 26, 27, 28] estimations.
For isotropic membranes, the stiffness matrix can be expressed in terms of the Lamé parameters $\lambda$ and $\mu$
@@
$K^{\prime\prime}_{\textrm{b}}=\frac{d^{2}V_{b}}{dr^{2}}\Bigr{|}_{r=r_{b}}=2a_{\mathrm{b}}^{2}D_{\mathrm{b}}$ (22)
@@
$\tilde{K}=\frac{K^{\prime\prime}_{\textrm{b}}}{6K^{\prime\prime}_{\textrm{a}}}\,,$ (26)

### v2

# First-principles derived force field for h-BN monolayer nanostructures: Applications to sheets, nanotubes and nanotori

@@ Abstract
In this work, we present an empirical force field for hexagonal boron nitride (hBN) monolayers, derived via a bottom-up strategy from first principles calculations. We aim to deliver a simple analytical force field for boron nitride which is efficient and applicable to large-scale simulations, without compromising its accuracy. The force field is developed with the goal to analytically reproducing the potential energy contributions arising from planar bond-stretching and bond-angle-bending deformations, as well as from out-of-plane torsional deformations, as obtained from periodic density functional theory calculations. Analytical anharmonic potential energy functions were employed to describe and parameterize the force field through a fitting process. The potential is applied for estimating in-plane (stiffness matrix and elastic constants) and out-of-plane (bending and Gaussian rigidity) mechanical properties of monolayer hBNs using both molecular mechanics calculations and analytical formulas. Additionally, we determined the elastic properties of more complex nanostructures, including hBN nanotubes and nanotori. By combining a continuum model for bent hollow beams with the definition of bending rigidity, we validate the model predictions against an analytical expression for the bending rigidity of large-diameter hBN nanotubes, expressed as a function of their radius and the in-plane Young modulus.
@@ I Introduction
Hexagonal boron nitride (hBN) has sparked significant interest within the scientific and industrial communities due to its unique properties. Being a two-dimensional material with a wide indirect bandgap (5.955 eV)[1], excellent chemical and thermal stability,[2] exceptional thermal conductivity[3] and good dielectric and optical properties,[4, 5], it is highly versatile for numerous applications.[6, 2, 5, 7] In the field of electronics, hBN has been successfully employed in transistors, integrated circuits, and capacitors, leading to enhanced performance and improved energy conversion efficiency.[8, 9, 10] Moreover, its remarkable thermal management capabilities have been utilized in electronic devices and green LEDs, effectively dissipating heat.[11] The optical properties of hBN have also proven valuable in optoelectronics and photonics,[12, 13] where it can serve as a transparent electrode, UV light emitter,[14] and as UV-detector.[5] In addition, hBN has been considered for applications in photocatalytic carbon dioxide reduction.[15] Furthermore, hBN’s chemical stability and lubricating properties have contributed to its use in protective coatings and high-temperature lubrication.[16, 17, 18]
@@
The mechanical properties of hBN have been investigated in the literature via experiments,[23, 24] ab-initio calculations,[25, 26, 27, 28, 29] and atomistic molecular dynamics (MD) simulations using Tersoff-based force fields in most cases.[30, 31, 32, 33, 34] In this work, we develop a semi-empirical force field for hBN monolayers encompassing both in-plane (bond-stretching and bond-angle-bending) and out-of-plane (torsional) interactions. The force-field parameterization results from a bottom-up fitting procedure in terms of matching the potential energy of properly deformed hBN geometrical structures as derived from first-principles density functional theory (DFT) calculations.
The functional expressions of the proposed potential are simple yet effective, capable of efficiently addressing very large hBN sheets with large-scale molecular dynamics simulations. The resulting force field is significantly more efficient than previous Tersoff-based implementations[30, 31, 32, 33, 34]. It has been employed to investigate the structural relaxation and elastic properties of various hBN-based nanostructures, including hBN monolayers, nanotubes (NT), and hBN nanotori.
In the following, we provide a detailed description of the parametrization of the in-plane (Section II.1) and out-of-plane (Section II.2) terms of the force field through the fitting of the potential energy of deformed hBN monolayers obtained from ab-initio calculations. Subsequently, in Section III.1 we determine the elastic constants of single-layer (SL) hBNs at T = 0 K, such as the components of stiffness matrix and the corresponding Lamé parameters, Young modulus, Poisson ratio, and bulk modulus. The aforementioned parameters are estimated via molecular mechanics calculations and with analytical closed-form formulas.[35] In Section III.2, we illustrate the variation of the potential energy of hBN nanotubes as a function of their radius and chirality. In the limit of infinite radius, we extrapolate the bending rigidity of the corresponding hBN monolayers to provide comparisons against analytical expressions. Finally, in Section III.3 following a similar procedure for the energy of folded hBN NT in order to form closed nanotori, the bending rigidity of hBN nanotubes is estimated.
@@ II Force Field Derivation
We considered the minimal 2-atom unit cell to simulate the in-plane deformations with an $8\times 8\times 1$ Monkhorst pack sampling of the reciprocal space. Benchmark calculations demonstrate that a k-point mesh with $k\geq 8$ along the periodic directions yields converged forces for the systems under consideration. To determine the bond-stretching potential term, the unit cell lattice vectors were scaled uniformly (Fig. 1a). Regarding the bond-angle-bending term, the unit cell was deformed in such a way (Fig. 1b) that all bond lengths assumed rigid and only the angles were varied in a symmetric fashion (see Fig. 3 below).
For the out-of-plane deformations, we considered deformed BN nanoribbons that were bent around a particular middle axis. Three different cases were considered: (i) bending an 18 atom unit cell around a zig-zag (ZZ) middle axis with only boron atoms lying on that axis (Fig. 1c top, where the rotational zig-zag axis is shown by the green dotted line denoted as ZZ), (ii) similarly bending a nanoribbon, with a 18 atom unit cell as well, around a zig-zag middle axis passing strictly through nitrogen atoms, and (iii) bending a nanoribbon, with a 22 atom unit cell, around an armchair (AC) middle axis (Fig. 1c bottom, where the rotational armchair axis is shown by the green dotted line denoted as AC). For the out-of-plane deformations, at least an $1\times 24\times 1$ Monkhorst pack sampling of the reciprocal space was used. In all cases, a vacuum gap larger than 2 nm was maintained along the aperiodic directions to prevent spurious interactions.
@@ II.1 In-plane terms of the force field
Figure 1: Schematic illustration of the deformed hBN structures considered in the DFT calculations. In-plane deformations: (a) affine deformation with a uniform scaling along lattice vectors (Eq. (2)) and (b) deformation altering in-plane angles and maintaining bond lengths (Eq. (4)). Green arrows indicate the displacement direction of box corners. Out-of-plane deformations: (c) folding an hBN ribbon by an angle $\varphi$ around a rotation axis (green dotted line) oriented along the ZZ (top) or AC (bottom) direction passing through its plane. In all cases, red dashed lines depict the edges of the unit cell used in the calculation. In (c), the box length in the aperiodic direction is much larger than the ribbon width to prevent spurious interactions; consequently, the unit cell edges normal to this direction are not shown. Supercells with three-unit cell replications along the periodic directions are drawn for clarity. The blue transparent (opaque) lines illustrate the periodic boundaries of the supercells before (after) the deformation.
The in-plane potentials of hBN were determined by deforming a periodic primitive cell with the following (unperturbed) lattice vectors: $\mathbf{a}_{1}=(a,0)$ and $\mathbf{a}_{2}=(\frac{a}{2},\frac{a\sqrt{3}}{2})$, where $a=\|\mathbf{a}_{1}\|=\|\mathbf{a}_{2}\|=\sqrt{3}\;l_{\textrm{BN}}$ is the lattice constant and $l_{\textrm{BN}}$ the B-N bond length. In the reference configuration ($\mathbf{R}^{\textrm{ref}}$), where there is no residual stress, the bond length equals $l_{\textrm{BN}}=1.4518$  and the lattice constant is $a=2.5146$; the error for $l_{\mathrm{BN}}$ and $a$ is lower than the reported significant digits.
@@
with $\epsilon$ being the engineering strain and $I_{2}$ the identity matrix of size 2; see illustration in Fig. 1a. This operation scales the bond lengths uniformly and does not modify the in-plane angles. The same method has been applied in Ref. 41 for the case of graphene monolayers.
The bonded interactions of the force field were then parameterized through fitting of the DFT energy per bond results obtained from the uniform scaling deformations, shown in Fig. 2 by circles, with the Morse potential
@@
where $D_{\textrm{b}}$ is the potential well depth, $r_{\textrm{b}}$ the equilibrium bond distance (fixed parameter), and $\alpha_{\textrm{b}}$ the inverse width parameter of the potential. The fitting of the DFT data with Eq. (3) is shown by the continuous line in Fig. 2, yielding the following parameter values: $D_{\textrm{b}}=5.35\pm 0.01$ eV, $\alpha_{\textrm{b}}=1.897\pm 0.004$-1, and $r_{\textrm{b}}=1.4518$. The calculated equilibrium B-N distance leads to a lattice constant $a=2.5146$  in excellent agreement with literature values [23, 42, 32, 26, 27, 28].
@@
The bond-angle-bending term of the force field is obtained through alteration of the in-plane angles of a pristine hBN, by $-x$ or $x/2$ while maintaining the equilibrium value of $1.4518$ in all B-N bonds, as shown in Fig. 3. [41] In this case, four angles within each hBN ring increase by $x/2$, and two angles decrease by $x$. This operation deforms the reference configuration nonaffinely. First, we set the origin of the first (1) and second (2) sublattice (e.g., the pink and blue atoms in Fig. 1b, respectively) to
@@ II.2 Out-of-plane terms of the force field
Note that hBN exhibits mirror symmetry with respect to normal planes passing through neighboring B-N atoms; hence, a single deformation suffices for determining folding around the AC direction. The torsional energy of the AC-folded hBNs, $U\mathrm{{}_{t}^{(a)}}$ is shown in Fig. 6 by black circles. On the contrary, hBN is asymmetric with respect to normal planes passing through second neighboring identical (B or N) atoms. To capture this effect in the case of folding around the ZZ direction, the torsional energy was averaged from two separate folding deformations, where the rotational axis passes through second neighboring either B or N atoms (see Fig. 5 left). The corresponding torsional energies $U\mathrm{{}_{t}^{(z)}}$ for the two separate ZZ folding deformations are shown by red open symbols in Fig.6, while their average is represented by filled red circles.
Note that the total deformation energy obtained by DFT in these cases includes contributions not only from torsional terms but also from bond-angle-bending terms. Before fitting the parameters of the out-of-plane potentials one must remove the latter contributions, in order to isolate the dependence of the pure torsional terms on the folding angle $\mathit{\phi}$. Details of this fitting procedure can be found in Ref. 44.
Figure 6: Torsional energy of hBN nanoribbons folded along AC (filled black circles) or ZZ (red open symbols) rotation axes. In the latter case, the results are shown by empty diamonds when the ZZ axis passes through N atoms and by empty circles when the axis passes through B atoms, as it is illustrated in Fig. 5 (left). The average of these two situations (filled red circles) is considered in the fitting procedure regarding the rotation around ZZ folding axis. Solid lines indicate the model’s fittings using Eqs. (10) and (11) for nanoribbons folded along the AC and ZZ axes.
As it is customary, we define the torsional angle, $\omega$, for a quadruple of atoms $(ijkl)$ as the dihedral angle of the planes defined by $(ijk)$ and $(jkl)$ atom triplets. One can distinguish two kinds of such angles, termed cis and trans. As in ref. 44, we used the following analytic formulas for the out-of-plane distortions as functions of $\mathit{\omega}$:
@@ III.1 In-plane Elastic Constants of hBN Monolayers
where now pair indices are denoted by single numbers, $xx=1$, $yy=2$, and $xy=6$. In doing so, the strain is described by the vector
@@
To determine the elements of the stiffness matrix via MM computations we utilize three fundamental deformation modes (elementary deformations), each characterized by a single non-zero strain component in Eq. 15; hence, the components of the stiffness matrix are derived from the slope of the corresponding stress-strain plots in the linear regime for small strains.[68]
@@
respectively; note that $\epsilon_{1}=\xi$ and $\epsilon_{2}=\xi$. The third deformation mode $\bm{\epsilon}=(0,0,\epsilon_{6})$ entails varying the angle $\hat{\gamma}$ between the lattice vectors of an initially orthorhombic box (shearing) upon applying the linear transformer
@@
where $\epsilon_{6}=\tan(\hat{\gamma}\xi)\approx\hat{\gamma}\xi$. The slope of the corresponding stress-strain plots was evaluated with a central-difference scheme with perturbation $\xi=\pm 10^{-7}$. For additional details regarding the determination of the stiffness matrix with explicit deformation, the interested reader is referred to the article by Clavier et al;[68] see Section 2.1 therein. Details regarding the calculation of the stress tensor can be found in Ref. 69.
@@
[…]
Table 2:  The lattice parameter a, the in-plane and out-of-plane (bending rigidity $D_{\mathrm{sheet}}$ and Gaussian rigidity $D_{\mathrm{G,sheet}}$) elastic constants of hBN monolayers, calculated through MM simulations using the proposed force field, along with comparisons of corresponding experimental and theoretical (atomistic or ab initio methods) values reported in the literature1.
@@
$a$ (Å) 2.5146 2.504 [23], 2.505 [42] 2.505 [32, 70], 2.504 [34] 2.512 [26], 2.51 [27], 2.504 [28]
[…]
$\nu$ 0.275 0.208 [71, 24], 0.200 [71, 72] 0.297 [32], 0.17 [73, 74] 0.216 [29], 0.2176 [26]
0.16 [73, 71], 0.186 [34]
[…]
256 [71, 24], 237 [71, 72] 295 [73, 74], 280 [73, 71], 267 [34] 275.8 [27], 256.245 [28]
[…]
$D_{\textrm{sheet}}$ (eV) 1.56 - 0.56 [32], 0.86 [76], 0.89 [77] 1.29 [25], 0.95 [28]
$D_{\textrm{G,sheet}}$ (eV) $-3.23$
1 The characteristic thickness ($l_{\mathrm{t}}$) for converting 3D stress to 2D units, Eq. (16), has been considered 3.34 Å  in Ref. [32, 75], 3.3 Å  in Ref. [28], 3.33 Å  in Refs. [31, 24, 70], and $r_{\mathrm{N}}+r_{\mathrm{B}}=1.55$ $+1.92$ $=3.47$  in Ref. [29] ($r_{\mathrm{N}}$ and $r_{\mathrm{B}}$ are the van der Waals radii of N and B atoms, respectively).
The components of the stiffness matrix and the corresponding elastic constants are in good agreement with the values reported in the literature from MD and ab initio simulations. In addition, the lattice constant conforms with the reported experimental and theoretical estimates.
For isotropic 2D membranes, the stiffness matrix can be expressed in terms of the Lamé parameters $\lambda$ and $\mu$
@@
$K^{\prime\prime}_{\textrm{b}}=\frac{d^{2}V_{\mathrm{b}}}{dr^{2}}\Bigr{|}_{r=r_{\mathrm{b}}}=2\alpha_{\mathrm{b}}^{2}D_{\mathrm{b}}$ (22)
@@
$\tilde{K}=\frac{K^{\prime\prime}_{\textrm{b}}}{6K^{\prime\prime}_{\textrm{a}}}\,$ (26)